FAR-INFRARED SPECTRA AND TWO-DIMENSIONAL POTENTIAL ENERGY SURFACES FOR THE OUT-OF-PLANE RING VIBRATIONS OF TETRAHYDROFURAN-3-ONE IN ITS S0 AND S_{1}(n,π^{∗}) ELEC-TRONIC STATES

\underline{SOO-NO LEE}, NIKLAS MEINANDER, PAUL SAGEAR AND JAAN LAANE, Department of Chemistry, Texas A&M University, College Station, Texas 77843-3255; ,.

The far-infrared spectra of tetrahydrofuran-3-one show ring-bending (100 - 120 cm^{−1}), ring-twisting (227 - 237 cm^{−1}), difference (115 - 137 cm^{−1}), sum (320 - 350 cm^{−1}) and overtone (200 - 225 cm^{−1}) bands. A two-dimensional potential energy surface for the twisting and bending was determined and this has minima corresponding to twisted conformations. A potential energy surface for the S_{1}(n,π^{∗}) excited state, based on previous fluorescence excitation data, was also calculated.